OpenStudy (anonymous):

Write an equation in standard form of the line that passes through (0,3) (2,-3)

OpenStudy (anonymous):

Find the slope using slope formula \[(\frac{ x_1 + x_2 }{ 2 }, \frac{ y_1 + y_2 }{ 2 })\] Then put it in point slope form \[y-y _{1}=m(x-x _{1})\] Then in slope intercept form : y=mx+b Then change it into Standard form: Ax+By=C